OpenStudy (anonymous):

The slope of a line is 13 and the point (3, 19) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

y-19 = 13 ---- x-3 y-19=13(x-3) y-19=13x-39 y=13x-20